OpenStudy (anonymous):

please help The formula for degrees Celsius is: C 5/9 = (F − 32), where F stands for degrees Fahrenheit. Part 1: Solve the equation for F and show all steps. Part 2: Determine how many degrees Fahrenheit 25 degrees Celsius is.

OpenStudy (allank):

The question requires you to make F the subject of the equation. In other words, play around with the equation to have it like F=......

OpenStudy (anonymous):

that doesnt make sense cause f is in a equation

OpenStudy (anonymous):

nop that not right

OpenStudy (anonymous):

i still dont understnad

OpenStudy (anonymous):

wait C 5/9 = (F − 32) so c= 9/5(F − 32)

OpenStudy (anonymous):

thats what i ment

OpenStudy (anonymous):

1. C 5/9 = (F − 32) so c= 9/5(F − 32) c=9F/5 - 288/5 c=9F/5 -57.6 c+57.6=9F/5 (c+57.6)5/9=F

OpenStudy (anonymous):

\[C=(F-32)/1.8\] \[1.8C = F-32\] \[F=1.8C+32\] Plug in C and solve for F

OpenStudy (anonymous):

1) multiply thru by 1.8 to get rid of it on the right side of the equation 2) Add 32 to both sides to get F by itself 3) If you do the same thing to both sides of an equation the equation essentially doesn't change, it is just manipulated.